# Paired Course and Dorm Allocation

### v3

@@ Abstract
Consider a university administrator who must assign students to both courses and dorms and runs (random) serial dictatorship independently in each market. While serial dictatorship is efficient for each individual matching problem holding the other fixed, Pareto improvements can be found via students jointly trading their allocated course and dorm. To align priorities with students’ relative preferences between courses and dorms, we introduce paired serial dictatorship: a novel mechanism where students signal relative preferences to influence their priority in each market. Any deterministic allocation that arises in equilibrium is Pareto efficient, pointing towards ad-hoc tie-breaking as the key barrier to optimality. When students differ only in relative preferences, paired serial dictatorship ex-ante Pareto dominates running random serial dictatorship in each market. Such gains are realized despite students never reporting full preferences, a key advantage over other combinatorial allocation mechanisms that scale poorly with problem size.
Keywords: Combinatorial Assignment, Course Allocation, Multidimensional Allocation.
@@ 1 Introduction
The courses a student takes and the dorm they live in are foundational aspects of any college experience. Unfortunately, course and dorm selection are difficult and stressful processes for many students, who often are unable to get their favorite schedule or residence. Universities also face constraints in the mechanisms they can use to screen students: In a standard mechanism design framework, the designer can leverage transfers to elicit information about preferences, but payments in this setting would be repugnant and unfair for institutional reasons. In particular, it would disadvantage students from low-income backgrounds, contrary to one of higher education’s goals of increasing economic mobility. As a result, many universities have adopted serial-dictatorship-like mechanisms, where:
  1. 1.
    Each student is first assigned a course enrollment time and a dorm selection time independently at random;
  2. 2.
    When it is their time to choose in either market, students can pick from any remaining unfilled courses or dorms.
In practice, priorities are often determined by seniority (so students who have been matriculated for longer have earlier times) with random tie-breaking between students of the same year, done independently across markets. This process, which we will refer to as independent RSD, ignores relative preferences between courses and dorms: A student who cares significantly more about their courses may receive a late course enrollment time but an early dorm selection time, while a student who cares significantly more about their dorm may receive an early course enrollment time but a late dorm selection time. As such, while serial dictatorship is efficient within each individual market, it may not be Pareto efficient when we instead consider the joint allocation of courses and dorms.
One natural solution to recover Pareto efficiency would be to randomly order students once and let them pick both their courses and dorm. However, this process generates the same ex-ante marginal distributions of courses and dorms allocated to each student as running random serial dictatorship independently. Furthermore, students face greater uncertainty and fairness is sacrificed: Ordering students only once leads to strong correlation between better courses and better dorms. As such, student satisfaction is much more sensitive to randomization when a singular random priority determines both their courses and dorm. Additionally, one student may receive both a better course schedule and a better dorm than another student, violating envy-freeness.
Beyond the issue of relative preferences, even fully eliciting and reporting preferences over bundles of courses is a difficult matter. Some courses may be pre- or co-requisites of other courses, making them complements. On the other hand, two different classes in different departments may cover similar material (Stanford has at least seven different introduction to probability theory courses; many of them can explicitly be used to replace one another), making them substitutes. Students often prefer a balance of different types of courses (for example, an ideal schedule may include two courses with problem sets and exams, one project-based course, and one course for fun that is more creative), [^2] Some students might even have the opposite preferences and prefer to only have exam-based courses or project-based courses. which can make courses complements or substitutes depending on the student’s other courses.
To illustrate the complexity of this issue, consider the problem of a Stanford economics student choosing courses for the 2024 Fall Quarter. Stanford offered 18 undergraduate economics courses that quarter. Removing the “core” classes which students all need to take and have similar preferences for, there are still 14 “elective” or “field” classes. [^3] This is also ignoring classes in other departments like statistics or computer science that can also count towards an economics major. Suppose this student would like to pick three economics classes as part of their overall course schedule. To fully express preferences, they would need to specify how they rank (${14\choose 3}=364$ potential alternatives. Even in this restricted environment, reporting is infeasible. A heuristics-based approach is also undesirable as there are often exceptions. For example, one attempt to reduce the problem size is to rule out any bundles that have classes that overlap. However, students often intentionally skip classes that have recorded lectures or well-written lecture notes, rendering that heuristic unusable. As such, serial dictatorship-like mechanisms are preferable: By asking students to choose from an available set of classes, the mechanism never needs to know any student’s full preference ordering and students never need to report or even explicitly compute that information.
With these considerations in mind, this paper considers a new mechanism for allocation of multiple goods without transfers, which we call paired serial dictatorship. Paired serial dictatorship works by asking students about their relative preferences between courses and dorms (through questions along the lines of, “On a scale of one to ten, how much do you value your courses compared to your dorm, with ten indicating only caring about your courses and one indicating only caring about your dorm?”). Using that report, students who report a higher relative preference for courses receive an earlier course enrollment time but a later dorm selection time, and vice versa for students who report a lower relative preference for courses. After orderings are determined, serial dictatorship is run in either market. This paper is the first to analyze how agents should be ordered under (random) serial dictatorship, in the case where there are two markets that can be interconnected.
Additionally, paired serial dictatorship is a minimal change over independently running random serial dictatorship. Current systems that run serial dictatorship for courses and dorms stays the same; the only difference is central school admin simply telling departments (e.g. Residential & Dining Enterprises for dorms and the Registrar’s Office for courses) what orderings to use instead of them drawing priorities randomly.
@@
We find that paired serial dictatorship induces the desired screening: Regardless of other students’ reports and preferences, students that care more about courses (or dorms) do indeed report that. This separation in turn leads to paired serial dictatorship allocating better courses to students that care more about courses and better dorms to students that care more about dorms. Whenever signaling and redistribution of priorities fully resolves coordination frictions—so randomized tie-breaking is immaterial to the resulting deterministic allocation—we arrive at Pareto efficient outcomes (Theorem 1). [^4] This result is similar in spirit to the analysis done in Feigenbaum et al. (2020), which considers a two-stage matching process for students to schools, incorporating preference updating while on waitlists. Under their mechanism, lotteries are reversed between the initial and second rounds of matching. Their key technical assumption is an “order condition” which guarantees that if one school is more popular than another in the initial round, it remains more popular in the second round. This assumption leads to the time schools reach capacity being well-behaved. In our model, the assumption of a deterministic allocation plays the same role, ensuring that the time each course or dorm reaches capacity is well-defined. While any student taking up a spot in a course or dorm has an externality on every other student, each student’s ability to optimize the signal they choose to reveal in equilibrium outweighs those externalities. As such, the need for ad-hoc tie-breaking when student reports are not sufficiently nuanced is the key driver of inefficiency, emphasizing how full randomization under independent RSD is the driver of suboptimality.
In the particular case where students have common orderings over courses or dorms and only differ in their relative preferences, paired serial dictatorship is an ex-ante improvement over independent RSD as each student can always “mimic” the latter mechanism (Proposition 6). If we ignore the combinatorial nature of course allocation and suppose each student only chooses one course, paired serial dictatorship with a given signal space yields the highest ex-ante aggregate utility among all mechanisms with that same signal space (Theorem 2). In many other mechanism design settings (including those where transfers are permissible), it is infeasible for the designer to fully elicit an agent’s type and may have to settle for some finite message space. Interpreting this result in the context of one of the goods being numeraire provides one way of overcoming such barriers in those settings.
@@
The remainder of the paper is organized as follows. We will first work through a simple example demonstrating the shortcomings of independent RSD and how paired serial dictatorship resolves inefficiencies. Next, Section 2 reviews related literature and discusses potential drawbacks of alternative course allocation mechanisms. Section 3 develops the model and provides two comparative statics about student behavior under paired serial dictatorship, while Section 4 analyzes welfare under paired serial dictatorship. In Section 5, an economy of students choosing courses and dorms is simulated to examine the performance of paired serial dictatorship when the assumptions in this paper do not hold. Finally, Section 6 concludes. Appendix A: Mechanism Details and Equilibrium collects notation used throughout the appendices, proves that paired serial dictatorship is well-defined, and shows that a pure-strategy Nash equilibrium exists. All proofs are collected in Appendix B: Omitted Proofs.
@@ 1.1 A Motivating Example
We begin with an example illustrating the potential issues with independent RSD. Suppose there are two students Chris ($C$) and Doug ($D$), two courses $c_{1}$ and $c_{0}$, and two dorms $d_{1}$ and $d_{0}$. Each student would like one course and one dorm, and each course or dorm has a capacity of one. Additionally, both students prefer $c_{1}\succ c_{0}$ and $d_{1}\succ d_{0}$ (i.e. subscripts denote quality), but Chris cares more about courses and Doug cares more about dorms. To parameterize this, let
@@
    If Chris chooses first in both markets, they receive $c_{1},d_{1}$ and a utility of $11$ while Doug receives $c_{0},d_{0}$ and a utility of 00.
@@
    If Chris chooses second in both markets, they receive $c_{0},d_{0}$ and a utility of 00 while Doug receives $c_{1},d_{1}$ and a utility of $11$.
@@
On the other hand, what happens if paired serial dictatorship is used? Suppose the designer simply asked students if they care more about courses or dorms. If both students give the same answer, then we return to uniform randomization of independent RSD, leading to each student receiving an (expected) utility of $(11+10+1+0)/4=5.5$. Otherwise, each student picks first in the market they say they care more about. As such, this mechanism induces the following normal form game:
@@
2 Literature Review
@@
This paper most directly contributes to the literature on course allocation. Contrary to other examples of one-sided matching or mechanism design without transfers, course allocation imposes additional structure: Students have preferences over bundles of courses that may exhibit both complementarities or substitutabilities while scheduling or balance constraints may rule out certain bundles altogether. The literature on course allocation, starting with Budish et al. (2010) and Budish (2011), has focused on the (approximately) competitive equilibrium from equal incomes (CEEI) mechanism, where students are each endowed with (approximately) the same amount of some fake currency, submit cardinal preferences over bundles of courses, and receive the market-clearing allocation. The CEEI mechanism has several attractive theoretical properties: Outcomes are approximately efficient, the mechanism is strategyproof when the number of students participating in the market is large, and each student gets at least their maximin share of utility and only ever envy another student’s bundle of courses up to a single course. Budish et al. (2017) details implementing the approximate CEEI mechanism at Wharton Business School where 1,700 students are assigned to 350 courses. Outcomes in practice are fair and efficient, reflected in student surveys where respondents report higher satisfaction and gains in perceived fairness. More recently, Kornbluth and Kushnir (2021) incorporates priorities (for example, students with seniority are favored over newer students) into the CEEI mechanism to form a new mechanism called Pseudo-Market with Priorities (PMP). They then use data from around 6,000 students across 7 colleges selecting from 756 courses to calibrate student utilities and evaluate PMP against deferred acceptance and serial dictatorship. PMP generally performs the best, leading to higher average utilities and lower standard deviations in utility with respect to random initializations. Unfortunately, such market-based mechanisms are a poor fit to addressing undergraduate course allocation writ large. In Kornbluth and Kushnir (2021), each college had no more than 1,642 students or 269 courses. Each year, Stanford has approximately 8,000 matriculated students and offers around 2,000 undergraduate courses per quarter; many other institutions are just as large, if not larger. It would be infeasible to ask students to provide cardinal utilities for every course they might be interested in, and solving for the market-clearing allocation would be difficult. Furthermore, joint trades of courses and dorms can still be found even when courses are allocated via CEEI.
@@
Instead, Budish and Cantillon (2012) considers the Harvard Business School “Draft” mechanism, where students are first assigned some order, choose their first class in order, choose their next class in reverse order, and so on. There are some similarities between student behavior under the Draft mechanism and paired serial dictatorship, formalized in the discussion following Proposition 2. Bichler and Merting (2021) investigates how randomized scheduling mechanisms compares to a first-come-first-serve mechanism. In particular, they find that the probabilistic serial mechanism generally outperforms random serial dictatorship when both are adapted to handle bundles of objects, but differences are small. Unfortunately, bundled probabilistic serial is not strategyproof, which students take advantage of in experiments using the mechanism. Bhalgat et al. (2011) carries out a similar comparison between probabilistic serial and random serial dictatorship, but for the case of one-to-one matching. Romero-Medina and Triossi (2024) evaluates deferred acceptance, immediate acceptance, and conditional acceptance algorithms in the context of course allocation.
@@
The other direction prior work has considered is extensions of the Shapley–Scarf economy to allow for multiple types of goods. Unfortunately, most results have been negative. Konishi et al. (2001) finds that even when preferences are strict and separable, there may not be any allocations in the (strict) core; if non-empty, the core may not be unique; and the set of allocations in the core may differ from the set of allocations that arise in competitive equilibrium. “Furthermore, there is no Pareto efficient, individually rational, and strategy-proof social choice rule.” Pápai (2001) analyzes what allocation rules are strategryproof, non-bossy (and satisfy citizen sovereignty, which states that for every possible allocation, there is some profile of preferences that leads to it). Only sequential dictatorships—allocation rules similar to serial dictatorships, but the order agents choose is endogenously determined by their reports—satisfy these three conditions. To get around the tension between Pareto efficiency and strategyproofness, Klaus (2007) considers second-best incentive compatible—a mechanism is second-best incentive compatible if no other incentive compatible mechanism Pareto dominates it—and shows that finding core allocations in the market for each type of good (independent of markets for other types of goods) is second-best incentive compatible. Anno and Kurino (2016) generalizes this result, finding that using strategyproof and non-wasteful allocation rules market-by-market results in a second-best incentive compatible allocation rule in the aggregate. We contribute to this literature by considering what can be attained when strategyproofness is dropped. In particular, the universe of reports available to each student is much smaller than the set of possible student types (a potential desiderata for allocation problems when reporting entire types is infeasible), which in and of itself leads to traditional notions of strategyproofness to break down. The Revelation Principle implies that there is a direct incentive compatible way to implement paired serial dictatorship; in the context of Anno and Kurino (2016), this means that paired serial dictatorship cannot Pareto dominate running each market independently. However, their results focus on deterministic rules and our results suggest that there are still large gains in ex-ante efficiency.
@@
Finally, we relate to the literature on mechanism design without transfers. Under a standard mechanism design framework, agent utilities are assumed to be quasilinear in money, leaving the designer with great flexibility to alter incentives by changing payments associated with different outcomes. Without transfers and a one-dimensional setting, we end up back at the one-sided matching world. When there are two dimensions like the setting of joint course and dorm assignment, the designer has more flexibility to use goods in one market as a shadow price on goods in the other market. Taking advantage of this, Kesten and Selcuk (2021) studies mediation in cases where two mediators have diametrically opposed ordinal preferences over two separate issues. A strategyproof and efficient mechanism exists only if every outcome in one dimension can be compensated by a sufficiently large concession in the other dimension, giving the designer sufficient flexibility to alter incentives. However, the literature on mechanism design without transfers has generally focused on the case where there are a large number of issues. The early paper of Jackson and Sonnenschein (2007) notes that any ex-ante efficient distribution over outcomes is attainable when enough copies of a decision are linked with itself. Ball and Kattwinkel (2023) works in the same setting but considers finite independent copies of a social choice problem and analyzes the performance of quota mechanisms, mechanisms which fix the distribution over outcomes across decisions. Horner and Guo (2015) and Balseiro et al. (2019) consider dynamic mechanism design without transfers in an infinite-horizon case and find that optimal mechanisms take the form of giving each agent a “utility promise” in the future which acts like a transfer. When there are enough periods, the designer can use design utility promises to act as transfers in the current period. Elokda et al. (2023) studies a similar problem but proposes a “karma” mechanism: Instead of utilizing utility promises, each agent has some amount of a fake currency called “karma” that they can either spend in the current period, or save and use in the future. The mechanism considered in this paper has similar characteristics: The signal a student sends to the mechanism about how much they care about their courses can be thought of as how much karma they’re willing to spend on an earlier course enrollment time, while any leftover karma goes towards obtaining an earlier dorm selection time.

### v4

# Simple Paired Combinatorial Assignment

@@ Abstract
Consider a university assigning students to courses and dorms. While many mechanisms are available, they each have their own drawbacks. Running serial dictatorship once for all goods is highly unfair, but running serial dictatorship separately for each matching problem is inefficient—Pareto improvements can be found via students jointly trading their allocated course and dorm. Alternatively, competitive equilibrium from equal incomes scales combinatorially in the number of items, making implementation and preference elicitation difficult. This paper considers paired serial dictatorship: a novel mechanism where agents signal relative preferences that determine their priority in each market. Any deterministic allocation that arises in equilibrium is Pareto efficient and envy-free, highlighting how seemingly innocuous tie-breaking is the key barrier to optimality and fairness. When agents differ only in relative preferences, paired serial dictatorship ex-ante Pareto dominates running random serial dictatorship independently in each market. Such gains exist even when agents behave simplistically.
Keywords: Combinatorial Allocation, Serial Dictatorship, Course Allocation.
@@ 1 Introduction
Combinatorial allocation problems are canonically difficult, especially when preferences exhibit complementarities and transfers cannot be used. While a wide range of potential mechanisms have been developed to solve this problem, each possesses its own benefits and drawbacks. One class of mechanisms involves adaptations of competitive equilibrium from equal incomes (CEEI) where each agent is endowed with some fictitious currency, reports valuations over bundles, and a competitive equilibrium of the resulting economy is found and implemented. Unfortunately, this approach faces two shortcomings. First is the non-existence of competitive equilibrium in the presence of complements, which has required modifications of the mechanism to settle for approximate competitive equilibrium from approximately equal incomes (see Budish (2011)). [^2] Nguyen and Vohra (2024) considers different approximations under alternative restrictions on preferences. The other major drawback is that agents need to report a cardinal utility for every potential bundle of goods to run CEEI, a problem that grows combinatorially in the number of goods. As such, the mechanism has only been used to allocate courses at small schools with limits on what preferences students can submit. [^3] In particular, students are only allowed to report a cardinal utility for each course and a (potentially negative) modifier for pairs of courses to indicate that they are either substitutes or complements. Students cannot report preferences over bundles of size three or greater (Budish and Cantillon, 2012). However, the mechanism is able to deliver (approximately) fair and efficient outcomes.
Another commonly used mechanism is random serial dictatorship (RSD), where agents are ordered in some way and choose their favorite bundle in sequence from what is still available. While this mechanism is Pareto efficient and simple for agents to engage with, it is highly ex-post unfair: Agents with a higher priority get to choose all their goods before agents with a lower priority get to choose any of their goods. On the other hand, if multiple instances of RSD were used to allocate different classes of goods in an attempt to restore fairness, ex-post Pareto efficiency is lost as running things independently (which will be called independent RSD) ignores agents’ relative preferences between different classes of goods.
Are there any mechanisms that achieve both the fairness of CEEI and the simplicity of serial dictatorship (while still guaranteeing efficiency)? This paper introduces and evaluates paired serial dictatorship—a novel mechanism for combinatorial allocation problems that can be decomposed into two components. In the first component, it must be that agents have single-unit demand and do not exhibit complementarities or substitutabilities with the remainder of the allocation problem. In the other component, arbitrary combinatorial preferences over bundles of any size are allowed. Paired serial dictatorship operates by first asking agents to report their relative preferences between the two components (e.g. “On a scale of one to ten, how much do you care about your courses versus your dorm?” in the context of course and dorm assignment) and uses those reports to order agents and run serial dictatorship in either market. Agents that report a higher relative preference in one component choose earlier in that component but later in the other component. This paper is the first to analyze how agents should be ordered under serial dictatorship, in the case where there are two classes of goods to be allocated. [^4] Additionally, paired serial dictatorship is a minimal change over independent RSD. Current systems that run serial dictatorship for courses and dorms stay the same; the only difference is central school administration simply telling departments what orderings to use instead of them drawing priorities randomly.
Our restriction of having one dimension of preferences being separable can arise by either making assumptions within a given allocation problem or by expanding the scope of the problem itself. This property also arises naturally in many of the canonical examples (taken from Budish (2011)) of combinatorial allocation without transfers:
  • In course allocation, certain fun or extracurricular courses (social dance at Stanford, Taylor Swift at Harvard, tree climbing at Cornell) are often orthogonal to a student’s primary coursework. Alternatively, universities may consider jointly allocating courses and dorms (which are often independently allocated via random serial dictatorship).
  • In worker scheduling, weekly tasks can be paired with the allocation of parking spaces or who has priority to take paid time off around holidays.
  • In the allocation of players to sports teams, goalies (or other positions) may exhibit less complementarities or substitutabilities with the rest of the team compared to other players. Alternatively, sports governing bodies may jointly allocate players and convenient travel schedules. [^5] Major league soccer (MLS) organizes a “SuperDraft” every year that runs serial dictatorship for new players that just graduated college. Teams are ordered by season performance, but ties are broken randomly among teams that performed the same. Furthermore, in 2018 the team that traveled the most traveled more than twice the distance of the team that traveled the least.
To fix ideas, we will use the application of allocating courses and dorms to students for the remainder of the paper: agents will be called students, the class of goods where agents have single-unit demand and are evaluated independently will be dorms, and the remainder of the market where combinatorial preferences are permitted will be courses.
@@
Paired serial dictatorship induces the desired screening: Students who care more about courses (or dorms) do indeed report that, regardless of other students’ reports and preferences. This separation, in turn, leads to paired serial dictatorship allocating better courses to students that care more about courses and better dorms to students that care more about dorms. However, multiple students will generally report the same relative preference, which means that tie-breaking is still required at times. Whenever this tie-breaking does not influence the resulting allocation—so signaling and redistribution of priorities fully resolves coordination frictions among students—the resulting deterministic allocation is Pareto efficient and ex-post envy-free (Theorem 1). As such, a significantly simpler mechanism is able to deliver the promises of CEEI without ever requiring students to report complete cardinal preferences over all potential bundles. [^6] This result is similar in spirit to the analysis done in Feigenbaum et al. (2020), which considers a two-stage matching process for students to schools, incorporating preference updating while on waitlists. Under their mechanism, lotteries are reversed between the initial and second rounds of matching. Their key technical assumption is an “order condition” which guarantees that if one school is more popular than another in the initial round, it remains more popular in the second round. This assumption is imposed to guarantee that the time each school reach capacity is well-behaved. In our model, the assumption of a deterministic allocation plays the same role, ensuring that the time each course or dorm reaches capacity corresponds precisely to the mass of students revealing a certain signal. While any student taking up a spot in a course or dorm has an externality on every other student, each student’s ability to optimize the signal they choose to reveal in equilibrium outweighs those externalities. As such, the need for ad-hoc tie-breaking when student reports are not sufficiently nuanced is the key driver of inefficiency as the mechanism simply does not have the necessary information to align priorities with preferences. In the worst case, full randomization under independent RSD is a key driver of suboptimality.
@@
The remainder of the paper is organized as follows. We first work through a simple example demonstrating the shortcomings of independent RSD and how paired serial dictatorship resolves inefficiencies. Next, Section 2 reviews related literature and discusses potential drawbacks of alternative course/combinatorial allocation mechanisms. Section 3 develops the model and provides two comparative statics about student behavior under paired serial dictatorship, while Section 4 analyzes welfare under paired serial dictatorship. In Section 5, an economy of students choosing courses and dorms is simulated to examine the performance of paired serial dictatorship when the assumptions needed for theoretical guarantees do not hold. Finally, Section 6 concludes. Appendix A: Mechanism Details and Equilibrium collects notation used throughout the appendices, proves that paired serial dictatorship is well-defined, and shows that a pure-strategy Nash equilibrium exists. All proofs are collected in Appendix B: Omitted Proofs.
@@ 1.1 A Motivating Example
We begin with an example illustrating the potential issues with independent RSD. Suppose there are two students, Chris ($C$) and Doug ($D$), two courses $c_{1}$ and $c_{0}$, and two dorms $d_{1}$ and $d_{0}$. Each student demands one course and one dorm, and each course or dorm has a capacity of one. Both students prefer $c_{1}\succ c_{0}$ and $d_{1}\succ d_{0}$ (i.e. subscripts denote quality), but Chris cares more about courses and Doug cares more about dorms. To parameterize this, let
@@
    If Chris chooses first in both markets, they receive $c_{1},d_{1}$ and a utility of $11$ while Doug receives $c_{0},d_{0}$ and a utility of 0.
@@
    If Chris chooses second in both markets, they receive $c_{0},d_{0}$ and a utility of 0 while Doug receives $c_{1},d_{1}$ and a utility of $11$.
@@
What happens if paired serial dictatorship is used? Suppose the designer simply asked students if they care more about courses or dorms. If both students give the same answer, we return to uniform randomization leading to each student receiving an (expected) utility of $(11+10+1+0)/4=5.5$. Otherwise, each student picks first in the market they say they care more about. As such, this mechanism induces the following normal form game:
@@
Furthermore, this result holds for arbitrary cardinal utilities of Chris and Doug (as long as Chris cares more about courses and Doug cares more about dorms); the mechanism arrives at the efficient outcome without ever knowing precise cardinal values. Unfortunately, the trade-off here is that a strictly dominant strategy may not always exist for more general economies. As such, paired serial dictatorship requires students to engage in equilibrium reasoning in exchange for simplicity when reporting preferences. In the course and dorm allocation problem (and many of the other problems discussed above), the allocation mechanism is run year after year and there is much discussion about what courses and dorms are in high demand. Students often talk about their draw times and what was available by the time they got to choose. Furthermore, students need not fully compute best responses to others’ underlying utility functions, but rather only need to compute expectations over when certain courses and dorms run out. Additionally, it is not the end of the world for any student even if they fail to behave optimally: Any courses and dorm that a student gets under paired serial dictatorship can also arise from independent RSD. Finally, simulations (in Section 5) suggest that the cost of bounded rationality is negligible.
2 Related Literature
@@
This paper most directly contributes to the literature on competitive equilibrium from equal incomes and its applications to course allocation. Contrary to other examples of one-sided matching or mechanism design without transfers, course allocation imposes additional structure: Students have preferences over bundles of courses that may exhibit both complementarities or substitutabilities while scheduling or balance constraints may rule out certain bundles altogether. The literature on course allocation, starting with Budish et al. (2010) and Budish (2011), has focused on the (approximately) competitive equilibrium from equal incomes (CEEI) mechanism, where students are each endowed with (approximately) the same amount of some fake currency, submit cardinal preferences over bundles of courses, and receive the market-clearing allocation. The CEEI mechanism has several attractive theoretical properties: Outcomes are approximately efficient, the mechanism is strategyproof when the number of students participating in the market is large, and each student gets at least their maximin share of utility and only ever envies another student’s bundle of courses up to a single course. Budish et al. (2017) details implementing the approximate CEEI mechanism at Wharton Business School where 1,700 students are assigned to 350 courses. Outcomes in practice are fair and efficient, reflected in student surveys where respondents report higher satisfaction and gains in perceived fairness. More recently, Kornbluth and Kushnir (2021) incorporates priorities (for example, students with seniority are favored over newer students) into the CEEI mechanism to form a new mechanism called Pseudo-Market with Priorities (PMP). They then use data from around 6,000 students across 7 colleges selecting from 756 courses to calibrate student utilities and evaluate PMP against deferred acceptance and serial dictatorship, clearing each college separately. PMP generally performs the best, leading to higher average utilities and lower standard deviations in utility with respect to random initializations. Unfortunately, such market-based mechanisms are a poor fit to addressing undergraduate course allocation writ large. In Kornbluth and Kushnir (2021), each college had no more than 1,642 students or 269 courses, but many other institutions have an order of magnitude more students or courses. It would be infeasible to ask students to provide cardinal utilities for every course they might be interested in, and solving for the market-clearing allocation would be difficult. Furthermore, joint trades of courses and dorms can still be found even when courses are allocated via CEEI.
@@
Instead, Budish and Cantillon (2012) considers the Harvard Business School “Draft” mechanism, where students are first assigned some order, choose their first class in order, choose their next class in reverse order, and so on. There are some similarities between student behavior under the Draft mechanism and paired serial dictatorship, formalized in the discussion following Proposition 2. Bichler and Merting (2021) investigates how randomized scheduling mechanisms compare to first-come-first-serve mechanisms. In particular, they find that the probabilistic serial mechanism generally outperforms random serial dictatorship when both are adapted to handle bundles of objects, but differences are small. Unfortunately, bundled probabilistic serial is not strategyproof, which students take advantage of in experiments using the mechanism. Bhalgat et al. (2011) carries out a similar comparison between probabilistic serial and random serial dictatorship, but for the case of one-to-one matching. Romero-Medina and Triossi (2024) evaluates deferred acceptance, immediate acceptance, and conditional acceptance algorithms in the context of course allocation.
@@
The other direction prior work has considered is extensions of the Shapley–Scarf economy to allow for multiple types of goods. Unfortunately, most results have been negative. Konishi et al. (2001) finds that even when preferences are strict and separable, there may not be any allocations in the (strict) core; if non-empty, the core may not be unique; and the set of allocations in the core may differ from the set of allocations that arise in competitive equilibrium. “Furthermore, there is no Pareto efficient, individually rational, and strategy-proof social choice rule.” Pápai (2001) analyzes what allocation rules are strategyproof, non-bossy (and satisfy citizen sovereignty, which states that for every possible allocation, there is some profile of preferences that leads to it). Only sequential dictatorships—allocation rules similar to serial dictatorships, but the order agents choose is endogenously determined by their reports—satisfy these three conditions. To get around the tension between Pareto efficiency and strategyproofness, Klaus (2007) considers second-best incentive compatible—a mechanism is second-best incentive compatible if no other incentive compatible mechanism Pareto dominates it—and shows that finding core allocations in the market for each type of good (independent of markets for other types of goods) is second-best incentive compatible. Anno and Kurino (2016) generalizes this result, finding that using strategyproof and non-wasteful allocation rules market-by-market results in a second-best incentive compatible allocation rule in the aggregate. We contribute to this literature by considering what can be attained when strategyproofness is dropped. In particular, the universe of reports available to each student is much smaller than the set of possible student types (a potential desiderata for allocation problems when reporting entire types is infeasible), which in and of itself leads to traditional notions of strategyproofness to break down.
@@
Finally, we relate to the literature on mechanism design without transfers. Under a standard mechanism design framework, agent utilities are assumed to be quasilinear in money, leaving the designer with great flexibility to alter incentives by changing payments associated with different outcomes. Without transfers and a one-dimensional setting, we end up back at the one-sided matching world. When there are two dimensions like the setting of joint course and dorm assignment, the designer has more flexibility to use goods in one market as a shadow price on goods in the other market. Taking advantage of this, Kesten and Selcuk (2021) studies mediation in cases where two mediators have diametrically opposed ordinal preferences over two separate issues. A strategyproof and efficient mechanism exists only if every outcome in one dimension can be compensated by a sufficiently large concession in the other dimension, giving the designer sufficient flexibility to alter incentives. However, the literature on mechanism design without transfers has generally focused on the case where there are a large number of issues. The early paper of Jackson and Sonnenschein (2007) notes that any ex-ante efficient distribution over outcomes is attainable when enough copies of a decision are linked with itself. Ball and Kattwinkel (2023) works in the same setting but considers finite independent copies of a social choice problem and analyzes the performance of quota mechanisms, mechanisms which fix the distribution over outcomes across decisions. Horner and Guo (2015) and Balseiro et al. (2019) consider dynamic mechanism design without transfers in an infinite-horizon case and find that optimal mechanisms take the form of giving each agent a “utility promise” in the future which acts like a transfer. When there are enough periods, the designer can use design utility promises to act as transfers in the current period. Elokda et al. (2023) studies a similar problem but proposes a “karma” mechanism: Instead of utilizing utility promises, each agent has some amount of a fake currency called “karma” that they can either spend in the current period or save and use in the future. The mechanism considered in this paper has similar characteristics: The signal a student sends to the mechanism about how much they care about their courses can be thought of as how much karma they’re willing to spend on an earlier course enrollment time, while any leftover karma goes towards obtaining an earlier dorm selection time.